# Conformal Calibration: Ensuring the Reliability of Black-Box AI in Wireless Systems

### v2

@@ II-B Reliable Prediction-Based Decision Making via Set Prediction
Given access to a prediction set $\Gamma$ with guaranteed coverage probability $1-\beta$, the optimizer can now replace the intractable average constraint $\mathbb{E}_{y}[R(x,y)]\leq\alpha$ with the known worst-case constraint $\max_{y\in\Gamma}R(x,y)\leq\gamma$. By construction, this guarantees that the KPI $R(x,y)$ remains below the threshold $\gamma$ with probability at least $1-\beta$. Through an appropriate selection of the parameters $\beta$ and $\gamma$, it can be shown that the original average constraint is satisfied [8, 7]. Indeed, recent results demonstrate that this approach is optimal for a broad class of optimization problems with statistical constraints under unknown distributions [9].
@@ II-D An Application: Power Control for Unlicensed Spectrum Access
Figure 3: True channel gain of the licensed user (dashed line), which is unknown and must be predicted after time zero, and (normalized) powers allocated by leveraging conformal prediction with two different confidence scores: a standard (negative) squared loss [8] and a multi-sample score introduced in [7] that yields multi-modal sets. Although both scores are based on the same AI model, a score that is better tailored to the problem at hand can yield significantly less conservative solutions (i.e., larger transmission powers).
@@ IV-A Robust and Reliable Prediction-Based Decision Making
The core principle behind online conformal prediction is to treat each new KPI observation as an opportunity to recalibrate. After each prediction is made and the KPI $R_{t}$ is measured, the basic threshold update rule works as follows: $\lambda_{t+1}=\lambda_{t}-\eta_{t}(R_{t}-\alpha)$, where $\eta_{t}$ is a step size. Thanks to this update, if the KPI $R_{t}$ – e.g., the latency level – happens to be higher than the target $\alpha$ at time $t$, the threshold $\lambda_{t+1}$ is lowered. This ensures that the prediction set will tend to increase in size (see Figure 2), yielding more conservative decisions at the next time step.
@@
Figure 5: Cumulative normalized SNR degradation with respect to the optimal beam as a function of time for a target normalized SNR degradation $\alpha=0.1$ (top figure), and average beam set size as a function of the target cumulative normalized SNR degradation (bottom figure). While both conventional and localized online conformal prediction meet the desired cumulative SNR degradation level as time goes on, the localized approach yields significantly smaller candidate beam set sizes, reducing training overhead.
@@ IV-B An Application: Downlink Beam Selection
Assume as the reliability measure $R_{t}$ the normalized level of SNR degradation as compared to the optimal beam within the codebook. For a degradation level $\alpha=0.1$, Figure 5 shows the cumulative average degradation over time (top figure) and the average candidate beam size as a function of the maximum tolerated SNR degradation (bottom figure) for online conformal prediction and localized online conformal prediction. As highlighted in the plot, both schemes can attain the desired time-averaged value of SNR degradation, with localized online conformal prediction yielding smaller candidate set sizes. This in turn entails the need to transmit fewer pilots.

### v3

@@ II-B Reliable Prediction-Based Decision Making via Set Prediction
Given access to a prediction set $\Gamma$ with guaranteed coverage probability $1-\beta$, the optimizer can now replace the intractable average constraint $\mathbb{E}_{y}[R(x,y)]\leq\alpha$ with the known worst-case constraint $\max_{y\in\Gamma}R(x,y)\leq\gamma$. By construction, this guarantees that the KPI $R(x,y)$ remains below the threshold $\gamma$ with probability at least $1-\beta$. Through an appropriate selection of the parameters $\beta$ and $\gamma$, it can be shown that the original average constraint is satisfied [7]. Indeed, recent results demonstrate that this approach is optimal for a broad class of optimization problems with statistical constraints under unknown distributions [8].
@@ II-D An Application: Power Control for Unlicensed Spectrum Access
Figure 3: True channel gain of the licensed user (dashed line), which is unknown and must be predicted after time zero, and (normalized) powers allocated by leveraging conformal prediction with two different confidence scores: a standard (negative) squared loss and a multi-sample score introduced in [7] that yields multi-modal sets. Although both scores are based on the same AI model, a score that is better tailored to the problem at hand can yield significantly less conservative solutions (i.e., larger transmission powers).
@@ IV-A Robust and Reliable Prediction-Based Decision Making
The core principle behind online conformal prediction is to treat each new KPI observation as an opportunity to recalibrate. After each prediction is made and the KPI $R_{t}$ is measured, the basic threshold update rule works as follows: $\lambda_{t+1}=\lambda_{t}-\eta_{t}(R_{t}-\alpha)$, where $\eta_{t}$ is a step size. Thanks to this update, if the KPI $R_{t}$ – e.g., the latency level – happens to be higher than the target $\alpha$ at time $t$, the threshold $\lambda_{t+1}$ is lowered. This ensures that the prediction set will tend to increase in size (see Figure 2), yielding more conservative decisions at the next time step. Assuming that the KPI of interest is decreasing in the size of the prediction set, this reactive behavior can be shown to ensure the satisfaction of the mentioned time-averaged deterministic constraints.
@@ IV-B An Application: Downlink Beam Selection
Figure 5: Cumulative normalized SNR degradation with respect to the optimal beam and corresponding beam set size as a function of time for a target normalized SNR degradation $\alpha=0.1$ (top two figures); and average beam set size, evaluated after $T=10000$ steps, as a function of the target cumulative normalized SNR degradation (bottom figure). While both conventional and localized online conformal prediction meet the desired cumulative SNR degradation level as time goes on, the localized approach yields significantly smaller candidate beam set sizes, reducing training overhead.
@@
Assume as the reliability measure $R_{t}$ the normalized level of SNR degradation as compared to the optimal beam within the codebook. This quantity may be periodically evaluated via full-codebook training, and the intermittent availability of feedback can be addressed as explained in [12]. For a degradation level $\alpha=0.1$, Figure 5 shows the cumulative average SNR degradation and the corresponding beam set size over time (top two figures) and the average candidate beam size as a function of the maximum tolerated SNR degradation (bottom figure) for online conformal prediction and localized online conformal prediction. As highlighted in the plot, both schemes can attain the desired time-averaged value of SNR degradation, with localized online conformal prediction yielding smaller candidate set sizes. This in turn entails the need to transmit fewer pilots.